# Randomized Approach to Matrix Completion: Applications in Collaborative Filtering and Image Inpainting

### v5

@@ 1.2 Outline of the paper
The rest of the paper is organized as follows: Section 2 covers related methods, our contribution, and the notations used. In Section 3, we present our method and provide formal analysis, particularly Theorem 3.2, which offers a lower bound on the probability of perfect reconstruction of the matrix, along with conditions on the minimum number of randomly selected columns and the minimum size of the sample of known elements of the reconstructed matrix that ensure achieving this estimate. Section 4 contains experimental results on synthetic and real datasets, confirming the effectiveness of the proposed method.
@@ 2.1 Low-rank matrix completion
The low-rank assumption is grounded in the observation that the real-world datasets inherently possess, or can be effectively approximated by, a low-rank matrix structure [10]. It plays a pivotal role in enhancing the efficiency of algorithms. The algorithms for the low-rank matrix completion typically leverage convex or non-convex optimization in Euclidean spaces [14, 9, 15, 16, 17, 18, 19], or employ Riemannian optimization [20, 21, 22]. The convex methods are the most mature and have strong and well-understood theoretical guarantees. The nuclear norm relaxation draws inspiration from the success of compressed sensing [10, 5, 23, 8] in which $\ell_{1}$ norm minimization is used to recover sparse signal instead of $\ell_{0}$ norm. The matrix rank may be expressed as the $\ell_{0}$ norm of the singular values vector $\mathbf{\sigma}$, while the nuclear norm is defined as its $\ell_{1}$ norm.
@@ 2.2 Column Subset Selection
Extensive research focuses on fully-observed matrices, which may not be practical for real-world applications [28, 29, 30, 31, 32, 33]. Bridging the research gap between CSS, CUR and matrix completion is a recently recognized challenge [35, 36, 37, 38]. In this context, two important questions emerge. The first one is about selecting meaningful columns (or rows) from partially observed data and applying CSS and CUR in matrix completion algorithms. Our method benefits from the fact that under standard assumptions about matrix incoherence, a high-quality column subset can be obtained by randomly sampling each column [39]. However, in the case of the coherent matrices, Wang and Singh [35] propose active sampling to dynamically adjust column selection during the sampling process, albeit with increased computational complexity.
@@ 3.1 Theoretical results
Figure 2: Coherence of the one-dimensional subspace $U\subset\mathbb{R}^{2}$. 2(a) depicts $U$ spanned by $(\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}})$ with the smallest coherence $\mu(U)=1$, while 2(c) presents $U$ spanned by one of the standard basis vectors $\mathbf{e}_{2}$ which achieves maximal coherence $\mu(U)=2$
@@ 4.2 Recommendation system
$\text{NMAE}=\frac{1}{|\Omega_{\text{test}}|(m_{\text{max}}-m_{\text{min}})}\sum_{(i,j)\in\Omega_{\text{test}}}|\mathbf{\hat{M}}-\mathbf{M}|,$ (14)
@@ Acknowledgments
Let $\tilde{U}_{\psi}$ denote subspace spanned by the $\psi$ first left singular vectors of $\mathbf{C}$ and let $P_{\tilde{U}_{\psi}}$ denote orthogonal projection on $\tilde{U}_{\psi}$. Then for parameter $\gamma>0$, with probability $1-2e^{-\gamma}$ we have
@@
$\|\mathbf{M}-\mathbf{\hat{M}}\|_{2}^{2}\leq\Delta+\frac{2n_{1}\tilde{r}\Delta}{\beta}.$ (26)
@@
Using the triangle inequality and the fact that Frobenius norm of a matrix is always greater than or equal to its spectral norm,
@@
$\leq\|\mathbf{M}-\mathbf{\tilde{U}}\mathbf{\tilde{U}^{T}}\mathbf{M}\|_{2}^{2}+\|\mathbf{\tilde{U}}\mathbf{\tilde{U}^{T}}\mathbf{M}-\mathbf{C}\mathbf{\hat{Z}}\|_{2}^{2}$
$\leq\|\mathbf{M}-P_{\tilde{U}}\mathbf{M}\|_{2}^{2}+\|\mathbf{\tilde{U}}\mathbf{\tilde{U}^{T}}\mathbf{M}-\mathbf{C}\mathbf{\hat{Z}}\|_{F}^{2}.$
@@
To bound the second component, we use the definition of the $\mathbf{Y}_{1}$ and $\mathbf{\hat{Y}}$, and the fact that the product of the Frobenius norms bounds Frobenius norm of the product of the two matrices.
$\|\mathbf{\tilde{U}}\mathbf{\tilde{U}}^{T}\mathbf{M}-\mathbf{\tilde{U}}\mathbf{\tilde{\Sigma}}\mathbf{\tilde{V}}^{T}\mathbf{\hat{Z}}\|_{F}^{2}$ $=\|\mathbf{\tilde{U}}\mathbf{Y}_{1}-\mathbf{\tilde{U}}\mathbf{\hat{Y}}\|_{F}^{2}$ (39)
$\leq\|\mathbf{\tilde{U}}\|_{F}^{2}\|\mathbf{Y}_{1}-\mathbf{\hat{Y}}\|_{F}^{2}$
Using (36) and the fact that $\|\mathbf{\tilde{U}}\|_{F}=\sqrt{\hat{r}}$, since columns of $\mathbf{\tilde{U}}\in\mathbb{R}^{n_{1}\times\tilde{r}}$ are orthonormal.
$\|\mathbf{\tilde{U}}\|_{F}^{2}\|\mathbf{Y}_{1}-\mathbf{\hat{Y}}\|_{F}^{2}\leq\frac{2n_{1}\tilde{r}\Delta}{\beta},$ (40)
implying
$\|\mathbf{M}-\mathbf{\hat{M}}\|_{F}^{2}\leq\Delta+\frac{2n_{1}\tilde{r}\Delta}{\beta}.$ (41)
@@
$\lambda_{\max}(\mathbf{H}^{i,j})$ $=\lambda_{\max}(\mathbf{\tilde{u}}_{i:}\mathbf{\tilde{u}}_{i:}^{T})$ (53)
$\leq\max_{1\leq j\leq r}|u_{ij}|^{2}$
$\leq\frac{\tilde{r}\mu(\tilde{U})}{n_{1}},$

### v6

@@ 1.2 Outline of the paper
The rest of the paper is organized as follows: Section 2 reviews related methods and presents our contribution. In Section 3, we present our method and provide formal analysis, particularly Theorem 3.2, which offers a lower bound on the probability of perfect reconstruction of the matrix, along with conditions on the minimum number of randomly selected columns and the minimum size of the sample of known elements of the reconstructed matrix that ensure achieving this estimate. Section 4 contains experimental results on synthetic and real datasets, confirming the effectiveness of the proposed method.
@@ 2.1 Low-rank matrix completion
The low-rank assumption is grounded in the observation that the real-world datasets inherently possess, or can be effectively approximated by, a low-rank matrix structure [10]. It plays a pivotal role in enhancing the efficiency of algorithms. The algorithms for the low-rank matrix completion typically leverage convex or non-convex optimization in Euclidean spaces [14, 9, 15, 16, 17, 18, 19], or employ Riemannian optimization [20, 21, 22, 23]. The convex methods are the most mature and have strong and well-understood theoretical guarantees. The nuclear norm relaxation draws inspiration from the success of compressed sensing [10, 5, 24, 8] in which $\ell_{1}$ norm minimization is used to recover sparse signal instead of $\ell_{0}$ norm. The matrix rank may be expressed as the $\ell_{0}$ norm of the singular values vector $\mathbf{\sigma}$, while the nuclear norm is defined as its $\ell_{1}$ norm.
@@ 2.2 Column Subset Selection
Extensive research focuses on fully-observed matrices, which may not be practical for real-world applications [29, 30, 31, 32, 33, 34]. Several works provide comprehensive overviews of CUR decomposition methods. For instance, [32] offers a detailed characterization of various forms of CUR, while [36] reviews different sampling strategies employed to construct CUR decompositions. Bridging the research gap between CSS, CUR and matrix completion is a recently recognized challenge [37, 38, 39, 40]. In this context, two important questions emerge. The first one is about selecting meaningful columns (or rows) from partially observed data and applying CSS and CUR in matrix completion algorithms. Our method benefits from the fact that under standard assumptions about matrix incoherence, a high-quality column subset can be obtained by randomly sampling each column [41]. However, in the case of the coherent matrices, Wang and Singh [37] propose active sampling to dynamically adjust column selection during the sampling process, albeit with increased computational complexity.
@@ 3.1 Theoretical results
Figure 2: Coherence of the one-dimensional subspace $U\subset\mathbb{R}^{2}$: Fig. 2(a) depicts $U$, spanned by $(\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}})$, with the smallest coherence $\mu(U)=1$. Fig. 2(b) illustrates the case where $21<\mu(U)<$ 2, while Fig. 2(c) presents $U$, spanned by one of the standard basis vectors $\mathbf{e}_{2}$, which achieves maximal coherence $\mu(U)=2$.
@@
Matrix Size Reduction in Stage I of CSMC
To ensure a high probability of success, we set $\gamma=\Omega(\ln r)$. Under the standard assumption of a low incoherence parameter $\mu_{0}(\mathbf{M})$, Theorem 3.2 guarantees exact recovery with $d=\mathcal{O}(r\ln r)$ fully observed columns. Consequently, instead of applying a memory-intensive convex solver to the full $n_{1}\times n_{2}$ matrix, it suffices to complete a much smaller matrix of size $n_{1}\times\mathcal{O}(r\ln r)$.
Sample Complexity of Stage II of CSMC
While the primary goal of the CSMC method is to provide an efficient two-stage matrix completion procedure, it is important to note that the second stage can also be viewed as a fast standalone matrix completion method under a specific observation model: a subset of $d$ columns of $\mathbf{M}\in\mathbb{R}^{n_{1}\times n_{2}}$ is fully observed, and a subset $\Omega$ of entries is observed uniformly at random. Under these assumptions, Theorem 3.2 guarantees exact recovery of $\mathbf{M}$, requiring $dn_{1}+|\Omega|$ observed entries for Stage II. Since $\tilde{r}\leq r$, $\mu_{0}(\mathbf{C})\leq\mu_{0}(\mathbf{M})$, and assuming a small incoherence parameter $\mu_{0}(\mathbf{M})$, the total number of observed entries simplifies to $\mathcal{O}\left(r(n_{1}+n_{2})\ln(n_{2}r)\right)$ for $\gamma=\Omega(\ln r)$. By introducing $n=n_{1}+n_{2}$, the sample complexity further simplifies to $\mathcal{O}\left(rn\ln(nr)\right)$. For comparison, nuclear norm minimization under uniform sampling for matrices with small incoherence parameters, requires $\mathcal{O}(nr\ln^{2}{n})$ observed entries, while recent modified non-convex gradient descent methods require $\mathcal{O}\left(r^{2}n\ln(n))\right)$ samples [44]. Column sampling approaches like CUR+ [41] achieve $\mathcal{O}\left(rn\ln(r)\right)$ complexity, and CCS requires $\mathcal{O}\left(r^{2}n\ln^{2}(n))\right)$ samples [3]. Thus, although the second stage of CSMC does not improve upon the best-known asymptotic sample complexity, it matches leading methods under standard incoherence assumptions and offers significant memory and computational efficiency advantages by solving an ordinary least squares problem.
Choice of sampling columns algorithm
To develop a fast and efficient matrix-completion method, we adopt a uniform column sampling strategy, which is computationally inexpensive and does not require access to all matrix entries. Moreover, Xu et al. [41] showed that under the standard incoherence assumption, uniformly sampled columns accurately approximate the column space of a matrix. We acknowledge that uniform sampling does not capture differences between columns. Alternative strategies, such as leverage score sampling [36, 30], volume sampling [45], or determinantal point process sampling [46], may yield better approximations but typically require full access to matrix entries. On the other hand, adaptive sampling methods, while potentially effective for matrices with high incoherence parameters, incur additional computational overhead [37, 39, 40].
@@ 4.2 Recommendation system
$\text{NMAE}=\frac{1}{|\Omega_{\text{test}}|(m_{\text{max}}-m_{\text{min}})}\sum_{(i,j)\in\Omega_{\text{test}}}|\hat{m}_{ij}-m_{ij}|,$ (14)
@@ Acknowledgments
Let $\tilde{U}_{\psi}$ denote subspace spanned by the $\psi$ first left singular vectors of $\mathbf{C}$ and let $P_{\tilde{U}_{\psi}}$ denote orthogonal projection onto $\tilde{U}_{\psi}$. Then for parameter $\gamma>0$, with probability $1-2e^{-\gamma}$ we have
@@
$\|\mathbf{M}-\mathbf{\hat{M}}\|_{2}^{2}\leq 2\Delta+\frac{4n_{1}\Delta}{\beta}.$ (26)
@@
By applying the triangle inequality, Cauchy-Schwarz inequality, and leveraging the fact that the Frobenius norm is an upper bound for the spectral norm,
@@
$\leq 2\|\mathbf{M}-\mathbf{\tilde{U}}\mathbf{\tilde{U}^{T}}\mathbf{M}\|_{2}^{2}+2\|\mathbf{\tilde{U}}\mathbf{\tilde{U}^{T}}\mathbf{M}-\mathbf{C}\mathbf{\hat{Z}}\|_{2}^{2}$
$\leq 2\|\mathbf{M}-P_{\tilde{U}}\mathbf{M}\|_{2}^{2}+2\|\mathbf{\tilde{U}}\mathbf{\tilde{U}^{T}}\mathbf{M}-\mathbf{C}\mathbf{\hat{Z}}\|_{F}^{2}.$
@@
To bound the second component, we use the definition of the $\mathbf{Y}_{1}$ and $\mathbf{\hat{Y}}$,
$\|\mathbf{\tilde{U}}\mathbf{\tilde{U}}^{T}\mathbf{M}-\mathbf{\tilde{U}}\mathbf{\tilde{\Sigma}}\mathbf{\tilde{V}}^{T}\mathbf{\hat{Z}}\|_{F}^{2}=\|\mathbf{\tilde{U}}(\mathbf{Y}_{1}-\mathbf{\hat{Y}})\|_{F}^{2}\leq\|\mathbf{\tilde{U}}\|_{2}^{2}\|\mathbf{Y}_{1}-\mathbf{\hat{Y}}\|_{F}^{2}=\|\mathbf{Y}_{1}-\mathbf{\hat{Y}}\|_{F}^{2}.$ (39)
Using (36), we obtain the following bound:
$\|\mathbf{M}-\mathbf{\hat{M}}\|_{2}^{2}\leq 2\Delta+\frac{4n_{1}\Delta}{\beta}.$ (40)
@@
$\lambda_{\max}(\mathbf{H}^{i,j})$ $=\lambda_{\max}(\mathbf{\tilde{u}}_{i:}\mathbf{\tilde{u}}_{i:}^{T})=\sum_{j=1}^{\tilde{r}}\tilde{u}_{ij}^{2}=\|\mathbf{\tilde{u}}_{i:}\|_{2}^{2}\leq\frac{\tilde{r}\mu(\tilde{U})}{n_{1}}$ (52)